OpenStudy (anonymous):

How would I do this?? A triangular playground has angles with measures in the ratio 8 : 7 : 5. What is the measure of the smallest angle? o 28 o 45 o 6 o 39

hartnn (hartnn):

ok, let the angles be 8x,7x,5x because they are in the ratio 8 : 7 : 5 also, sum of angles in a triangle is 180 so you have 8x+7x+5x=180 can u solve this to find x ?

OpenStudy (anonymous):

I have a question. So, to solve other problems like this one, would we always put it in that format? For instance, 5 : 7 would be 5x + 7x = 180 20x = 180 Divide both sides by 180 x = 9

hartnn (hartnn):

the triangle has 3 sides, so it would never be like 5:7 but for any other case , if the ratio is 5:7 u can safely take as 5x and 7x

OpenStudy (anonymous):

Okay I found x for the problem; I got 9. What do we do next?

hartnn (hartnn):

thats correct, x=9 and smallest angle is 5x =... ?

OpenStudy (anonymous):

Oh okay 5(9) = 45

OpenStudy (anonymous):

Thank you for your help! :) Would you mind helping me with one more?

hartnn (hartnn):

welcome ^_^ sure, ask.

OpenStudy (anonymous):

Write an equation in slope intercept form of the line through points P(6,7) with slope -2.

OpenStudy (anonymous):

I think it would be y = -2x + 7

hartnn (hartnn):

the point slope form is y-y1 = m(x-x1) to get slope -intercept form, you isolate y y-7=(-2)(x-6) now isolate y from this.

OpenStudy (anonymous):

y - 7 = -2x + 12 Add 7 on both sides y = -2x + 19

OpenStudy (anonymous):

Thanks again! :)

hartnn (hartnn):

yeah, thats correct

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Have a nice day/night! ^^

hartnn (hartnn):

you too :)

OpenStudy (anonymous):

Thank you for the medal! Much appreciated!